OpenStudy (anonymous):

Solve. lx+5l = 9 1. {-4,14} 2. {4,14} 3. {-14,-4} 4. {-14,4}

OpenStudy (alexandervonhumboldt2):

|x+5|=9 this is the absolute value: thus x+5=9 and -x-5=9 x+5=9 => x=4 -x-5=9 => x=-14 thu the roots are 4 and -14

OpenStudy (anonymous):

Thank you! :)